OpenStudy (anonymous):

find the x-intercepts of f(x)=4x^2-20

OpenStudy (anonymous):

x intercepts occur when y or f(x) = 0. so 0 = 4x^2 - 20 4x^2 = 20 x^2 = 5 \[x = \sqrt{5}\] or \[x = -\sqrt{5}\]

OpenStudy (anonymous):

That's the answer I got when solving 4x^2=20. So the x-intercepts are the same as the solution to the equation?

OpenStudy (anonymous):

the x-intercept is not the solution to the function f(x) = 4x^2 - 20. it's only the solution when f(x) = 0.

OpenStudy (anonymous):

Part a) " Solve 4x^2=20 " Part b) Find the x-intercepts of f(x)=4x^2-20 Part a) answer is the same answer you gave for the function

OpenStudy (anonymous):

Ahh the answer to part b) should be ( sqrt5,0) (-sqrt5, 0)